OpenStudy (anonymous):

What is the density of oxygen gas (O2) at 2.3 atmospheres and 305 Kelvin? Show all of the work used to solve this problem.

OpenStudy (anonymous):

There are two ways to solve this. The easier way involves knowing the density equation: D = (PM)/(RT) Where P is the pressure in atmospheres, M is the molar mass of the gas in grams per mole, R is the universal gas constant (0.0821 L*atm/mol*K), and T is the kelvin temperature. D = (2.3 atm)(32.00 g/mol) / (0.0821 L*atm/mol*K)(305 K) = 2.9 g/L The second way involves integrating your knowledge of density with the ideal gas law: You know that density is mass divided by volume: D = m / V The mass of the gas can be converted from the number of moles, if we could only know that. Fortunately we can work out the number of moles from the ideal gas law: PV = nRT Where n = the number of moles. Of course we don't know the volume of the gas, but that's okay...we don't have to. We can simply imagine a gas sample with a volume of 1.00 L. (2.3 atm)(1.00 L) = n(0.0821 L*atm/mol*K)(305 K) 2.3 atm*L = n(25.0 L*atm/mol) n = 0.092 mol So we need to convert 0.092 mol to grams: 0.092 mol O2 x (32.00 g O2 / 1 mol O2) = 2.9 g O2 Going back to the density formula: D = m / V = (2.9 g) / (1.00 L) = 2.9 g/L

OpenStudy (anonymous):

k thanks

OpenStudy (anonymous):

u r welcome